OpenStudy (anonymous):

evaluate the polynomial for x=2 3x^2-3x+7 I got 13 as my answer is this correct

OpenStudy (inkyvoyd):

Correct based off of my mental math.

Directrix (directrix):

3x^2-3x+7 3(2)^2 - 3 (2) + 7 = 3*4 - 6 + 7 =12 - 6 +7 = 13